### v3

# SuperMat: Physically Consistent PBR Material Estimation at Interactive Rates

Abstract
Decomposing physically-based materials from images into their constituent properties remains challenging, particularly when maintaining both computational efficiency and physical consistency. While recent diffusion-based approaches have shown promise, they face substantial computational overhead due to multiple denoising steps and separate models for different material properties. We present SuperMat, a single-step framework that achieves high-quality material decomposition with one-step inference. This enables end-to-end training with perceptual and re-render losses while decomposing albedo, metallic, and roughness maps at millisecond-scale speeds. We further extend our framework to 3D objects through a UV refinement network, enabling consistent material estimation across viewpoints while maintaining efficiency. Experiments demonstrate that SuperMat achieves state-of-the-art PBR material decomposition quality while reducing inference time from seconds to milliseconds per image, and completes PBR material estimation for 3D objects in approximately 3 seconds. The project page is at https://hyj542682306.github.io/SuperMat/.
Figure 1: We present an efficient method for high-quality material decomposition in both 2D and 3D. Left: Albedo and RM predictions of SuperMat on rendered images under unknown illumination. The first row shows inputs with uneven lighting, the second row displays albedo predictions, and the third row presents RM predictions. Middle: Rendering results of a 3D object with PBR materials obtained by our method under novel lighting. Right: Our method is efficient and produces high-quality results.
1 Introduction
Decomposing physically-based materials from images into their constituent properties (e.g., albedo, metallic, and roughness maps) remains a central challenge in computer vision and graphics. This task is essential for creating high-fidelity 3D assets across numerous applications, from film production and game development to virtual and augmented reality pipelines. While recent advances have shown promising results [5, 41, 23], the problem remains challenging due to its ill-posed nature, particularly given unknown lighting conditions and complex interactions between different material properties. Our work addresses two key objectives: 1) efficient decomposition of high-fidelity material maps from 2D images, and 2) extension of this capability to 3D objects, enabling accurate material estimation across complex geometries.
Previous works in this domain generally fall into two main categories: optimization-based and data-driven approaches. 1) Optimization-based methods [2, 43, 44, 15, 27, 16, 11, 21, 42, 14, 6] employ differentiable rendering and neural representations to reconstruct 3D objects with material parameters by minimizing the difference between rendered and target images. While these methods achieve reasonable view consistency, they often compromise material quality, producing artifacts such as blurred boundaries and physically implausible parameters, particularly in metallic and roughness properties. 2) Data-driven approaches [36, 47, 20], particularly recent ones [41, 5, 23], leverage large diffusion models and reframe material decomposition as a conditional image generation task. However, these methods tend to focus solely on individual material properties, overlooking the interactions between different types of materials and combined impact on rendered appearances. Even minor inaccuracies in individual material estimates can propagate through the rendering pipeline, leading to substantial visual artifacts. Moreover, current methods face considerable computational overhead - optimization-based approaches require costly geometric optimization, while diffusion-based methods demand multiple denoising steps (30-50) during inference, causing a critical bottleneck for real-time applications.
To address these challenges, we propose a deterministic one-step framework that achieves both efficient and high-quality material decomposition. First, inspired by HyperHuman [24], we introduce structural expert branches within the UNet architecture. These specialized branches enable concurrent processing of multiple material properties while maintaining domain-specific expertise, allowing our model to decompose albedo, metallic, and roughness maps simultaneously with minimal computational overhead. This design significantly improves consistency between different material properties while avoiding the linear cost scaling typically required when estimating multiple material attributes. Second, we address the computational inefficiency of traditional diffusion models by reformulating the DDIM sampling process [12, 37]. Through careful analysis and modification of the scheduler, we achieve single-step inference without compromising quality, dramatically reducing inference time. Most importantly, this single-step formulation enables end-to-end training of the diffusion model, allowing us to compute the perceptual and re-render loss on final predicted material textures, which enhances shadow removal and improves the joint performance of decomposed materials under varying lighting conditions. These advances culminate in SuperMat, a highly efficient image space material decomposition model that estimates high-fidelity albedo, metallic, and roughness maps at millisecond-scale speeds.
For practical 3D applications, we further extend SuperMat’s capabilities to handle 3D objects. We design a UV refinement network that processes SuperMat’s per-view predictions in UV space, addressing the challenge of maintaining consistency across different viewpoints while preserving the efficiency of our 2D approach. Integrated with SuperMat, this pipeline processes a textured 3D model in approximately 3 seconds, producing high-quality material maps that maintain consistency across arbitrary viewpoints and lighting conditions.
In summary, our contributions are:
  • We propose SuperMat, a novel image space material decomposition model that generates physically consistent albedo, metallic, and roughness maps simultaneously with a single-step inference process. Our method achieves state-of-the-art performance on diverse datasets, significantly outperforming previous diffusion-based works in speed.
  • With the help of structural expert branches and single-step inference, we successfully achieve end-to-end training with perceptual loss and re-render loss on diffusion models in the material decomposition process, significantly enhancing both material prediction quality and rendering fidelity.
  • We propose an efficient PBR material decomposition method for 3D models. Combining SuperMat and our UV refinement network, this pipeline generates high-quality PBR materials for 3D objects from baked textures in approximately 3 seconds.
2 Related Works
Material Decomposition. Recognizing the importance of material information in 3D assets, researchers have increasingly focused on material decomposition for objects and scenes. Several inverse rendering approaches [42, 14, 6] aim to reconstruct material-rich 3D representations from single or multi-view images under fixed lighting models. Although assumptions involving extra constraints can simplify the problem and yield valuable insights, they also limit the practical applicability of these methods. Other works [2, 43, 44, 15, 27, 16, 11, 21] focus on estimating both material properties and unknown illumination. As available information decreases, researchers have found that in addition to inverse control over material generation through rendered results, directly acquiring image space material priors is feasible. This shift has inspired a new wave of material decomposition methods [36, 47, 20, 41, 5, 23, 25, 39] targeting 2D images captured or rendered under unknown lighting conditions. However, challenges remain, as limitations in training data, network architecture, and training methods continue to leave room for substantial improvements in decomposition quality.
Diffusion Models. Recently, researchers have applied diffusion models to a range of tasks [29, 4, 35, 30, 45], demonstrating the effectiveness of casting outcome prediction as a conditional image generation process. By leveraging prior knowledge from large pre-trained models, recent studies [13, 34, 18, 9, 46, 10, 38, 5, 23] have achieved performance improvements over previous methods in tasks like depth, normal, and material estimation. However, incorporating large models introduces efficiency challenges, particularly high memory requirements and the time-intensive, multi-step denoising process needed for inference. Inspired by insights from recent work [24, 12, 22], our approach addresses these issues by significantly reducing memory and inference time, improving the feasibility of applying these models in real-world applications.
Generation in UV Space. UV maps are advantageous for fully encapsulating the appearance information of 3D models, leading several works to explore 3D object manipulation directly within the UV space. By representing the texture of an object as a learnable image, these methods [32, 3, 40] enable texture generation, inpainting, and editing, providing new avenues for 3D tasks and allowing 2D image-space techniques and insights to be applied to 3D appearances. Building on this approach, recent methods [1, 26] have applied diffusion models to refine UV maps with promising results. In this spirit, to extend SuperMat’s capability to 3D, we fine-tune a UV-space refinement diffusion model specifically.
3 Method
Figure 2: Overview of the SuperMat framework. The UNet architecture incorporates structural expert branches for albedo and RM estimation, enabling parallel material property prediction while sharing a common backbone. During training, we leverage a fixed DDIM scheduler and optimize the network end-to-end using both perceptual loss and re-render loss. The marked region shows the inference pipeline, where our model achieves single-step inference for concurrent material map generation. Input geometric information (normal, position) and environment maps are used to compute the re-render loss.
In this paper, we propose SuperMat, an image space material decomposition model fine-tuned from Stable Diffusion that efficiently generates high-quality albedo, metallic, and roughness maps simultaneously. Given a rendered or captured image of an object under an unknown lighting condition from a specific viewpoint, SuperMat generates the corresponding material maps for the model from that same viewpoint. As shown in Figure 2, SuperMat is a 2D image-to-image one-step model with structural expert branches that simultaneously generates both albedo and RM (Roughness, Metallic) maps, supporting single-step inference. Compared to previous diffusion-based methods, this model produces superior results in less time and with reduced memory usage. Our single-step formulation allows direct generation of material maps during training, in contrast to prior diffusion methods that predict noise ($\epsilon$) or velocity (v) terms. This enables end-to-end optimization with both perceptual loss on the generated maps and re-render loss from novel lighting conditions, leading to improved material quality and rendering fidelity.
While SuperMat achieves efficient single-image material decomposition, extending this capability to 3D objects presents additional challenges. Instead of relying on optimization-based approaches like score distillation sampling (SDS) that are computationally intensive and prone to mode collapse, we propose an efficient two-stage pipeline for material estimation on 3D objects. Given a 3D model with only RGB texture, we first apply SuperMat to decompose materials from multiple viewpoints. We then introduce a UV refinement network that consolidates these view-dependent predictions into complete, high-quality UV texture maps through single-step inference. As shown in Figure 3, this pipeline enables PBR material estimation for 3D objects in approximately 3 seconds per model.
Next, we first introduce the rendering model we use, followed by a detailed description of all the key components.
3.1 Physically Based Rendering (PBR) Model
We use the Cook-Torrance Bidirectional Reflectance Distribution Function (BRDF) [7] based on microfacet theory to define the materials and establish the rendering model. Following [4], for a point with coordinates $p\in\mathbb{R}^{3}$, albedo $a\in\mathbb{R}^{3}$, metallic $m\in\mathbb{R}$, roughness $r\in\mathbb{R}$, and normal variation $n\in\mathbb{R}^{3}$, the PBR result $L$ observed from viewpoint $c\in\mathbb{R}^{3}$ is given by:
$\begin{split}L(p,\omega)=a(1-m)\int_{\Omega}L_{i}(p,\omega_{i})(\omega_{i}\cdot n)\mathrm{d}\omega_{i}\\ +\int_{\Omega}\frac{DFG}{4(\omega\cdot n)(\omega_{i},n)}L_{i}(p,\omega_{i})(\omega_{i}\cdot n)\mathrm{d}\omega_{i},\end{split}$ (1)
where $\omega$ represents the direction of the outgoing light from point $p$ to $c$, i.e., the viewing direction. $L_{i}$ denotes the incident light from the direction $\omega_{i}$, and $\Omega=\{\omega_{i}:\omega_{i}\cdot n\geq 0\}$ represents the hemisphere of normals. $D$, $F$, and $G$ are the distribution, Fresnel, and geometry functions, respectively. For the integral part, the time complexity of Monte Carlo methods is unacceptable. Given that we use ambient lighting as the light source, we compute it efficiently using the split-sum method [27].
3.2 SuperMat - Image Space Material Decomposition
Previous works [41, 5, 23] have demonstrated that casting material estimation as a conditional latent diffusion process, leveraging knowledge from large pre-trained diffusion models like Stable Diffusion [33], is effective. However, for our task that requires simultaneous generation of multiple material maps, three key issues remain: 1) High model overhead: Fine-tuning Stable Diffusion with its original architecture can only estimate a single material type. Even when combining metallic and roughness into an RM map, treating it as a single material still requires two diffusion models, one for albedo and another for RM, to meet the rendering model’s requirements. This results in doubled training time and increased inference space overhead. 2) Slow inference speed: With the widely used DDIM scheduler, diffusion models require 30 to 50 denoising steps to produce satisfactory results, which can be inefficient and time-consuming. 3) Incomplete decomposition effect: Due to the high computational cost of backpropagation in multi-step inference, prior diffusion-based methods trained on predicted noise from single-step denoising rather than directly supervising the fully denoised material output. This limitation restricts the model’s ability to specialize in the decomposition task and impedes the application of more task-specific techniques.
Structural Expert Branches. Our solution to the first problem is to introduce a structural expert branch architecture [24]. Previous methods are limited to using a single diffusion model to estimate one material type. Although training separate models for different material maps is an intuitive approach, it introduces issues such as implementation complexity, high overhead, and inconsistencies between materials. Considering that, despite their different representations, materials fundamentally reflect an understanding of object features, we incorporate structural expert branches into the UNet model, allowing a single diffusion model to predict multiple material maps. Specifically, we design two expert branches, one for predicting albedo, and the other for predicting RM. For each expert branch, we replicate the UNet’s last UpBlock, and conv_out, while the rest of the UNet architecture remains shared between the branches. The modified model first extracts and transforms encoded image features through the shared modules. These features are then separated into 2 copies by duplicating before the last UpBlock, with each branch ultimately predicting results for its respective domain. The expert branches specialize in material-specific features, while the shared modules focus on commonalities. In this way, we only add a small number of parameters, achieving the same performance as previous methods that would otherwise require twice the scale.
Single-Step Inference and End-to-End Training. To address the second problem, we analyze the DDIM scheduler of diffusion models. Previous works have highlighted a flaw in the implementation of the DDIM scheduler [12, 22], which limits the potential of single-step inference in Stable Diffusion: In the default leading timestep setting of DDIM, for single-step prediction, the model is provided with a timestep ($t\text{=}1$) suggesting the input is essentially denoised, yet in reality, the input is pure noise. The discrepancy between the timestep and the actual noise level causes the single-step result to be flawed. To align the timestep with the noise level received by the model, we adjust the scheduler to use the trailing setting ($t\text{=}T$ for single-step), correcting this issue and enabling single-step inference, significantly reducing the time overhead.
Figure 3: Material Decomposition for 3D objects. Applying SuperMat’s powerful decomposition to 3D objects, we can obtain an object with high-quality materials in just three inference steps.
More importantly, after this correction, we can fine-tune the one-step model end-to-end, effectively resolving the third issue. For a diffusion model with the uncorrected timestep setting, fine-tuning the diffusion model end-to-end is impractical, because generating a clean material map typically requires multiple denoising steps, leading to prohibitive computational complexity due to the gradient calculations. As an alternative, previous methods [41, 5, 23] rely on using the noise latent obtained from a single denoising step of the UNet for loss calculation. However, this approach shifts the model’s training objective from material decomposition to general image denoising, which hinders specialization for this task. In contrast, our corrected inference pipeline can generate fully denoised material maps in a single denoising step, making end-to-end fine-tuning both feasible and efficient for material decomposition. Specifically, with the clean material estimations, we compute perceptual losses [17] between the predicted albedo map $k_{d}$, the RM map $k_{rm}$, and the ground truth maps $\hat{k}_{d}$, $\hat{k}_{rm}$:
$\begin{split}\mathcal{L}_{p}(\hat{y},y)=\sum_{j}\frac{1}{C_{j}H_{j}W_{j}}||\phi_{j}(\hat{y})-\phi_{j}(y)||_{1},\\ \mathcal{L}_{m}=\mathcal{L}_{p}(\hat{k}_{d},k_{d})+\mathcal{L}_{p}(\hat{k}_{rm},k_{rm}),\end{split}$ (2)
where $\phi$ is a 16-layer VGG network, and $\phi_{j}$ represents the output of the $j$-th layer, with a shape of $C_{j}\times H_{j}\times W_{j}$. In this paper, we use the $4$th, $8$th, $12$th, and $16$th layers to compute the feature reconstruction loss.
Note that, following the approach of previous work [12], we use “Zeros” as the noise type during fine-tuning, meaning no random noise is added.
Re-render Loss in DM Fine-tuning. Data-driven decomposition diffusion models typically focus on the distribution of individual material maps, yet they often overlook the interactions between different types of material maps. In other words, previous methods compute the loss between the predicted values and the ground truth without utilizing material information for further operations or calculations, which could provide stronger guidance for the model. As a result, although the predicted materials may appear close to the ground truth, the rendered output may not align with expectations. While this can be frustrating, it is understandable, as their training methods only produce meaningless predicted noise at each iteration. In contrast, our design, which incorporates structural expert branches and single-step prediction, allows us to obtain all denoised material maps during the training iterations. Using this information, we introduce the re-render loss, which involves rendering the model in a new lighting environment and comparing the result to the ground truth. Specifically, at each training iteration, we obtain $k_{d}$, $k_{rm}=(k_{r},k_{m})$, along with the true normal map $\hat{k}_{n}$, position map $\hat{k}_{p}$, camera viewpoint $\hat{c}$, and an environment light map. Then we use these information and the rendering model $\mathcal{R}$ described in Sec. 3.1 to compute the predicted material’s performance under new illumination. Similarly, with the groundtruth albedo $\hat{k}_{d}$ and RM $\hat{k}_{rm}$, we compute the ground truth of relighting. The perceptual loss between these two renderings defines our re-render loss:
$\mathcal{L}_{re}=\mathcal{L}_{p}(\mathcal{R}_{\hat{k}_{n},\hat{k}_{p},\hat{c}}(\hat{k}_{d},\hat{k}_{m},\hat{k}_{r}),\mathcal{R}_{\hat{k}_{n},\hat{k}_{p},\hat{c}}(k_{d},k_{m},k_{r})).$ (3)
3.3 UV Refinement Network
Albedo Metallic Roughness Relighting Time(s)↓
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Derender3D [36] 22.2606 0.8724 0.1622 - - - - - - - - - 0.16
IIR [47] 20.9665 0.8679 0.1629 16.8817 0.8118 0.2364 19.9573 0.8749 0.1578 21.5253 0.8901 0.1279 0.04
IID [20] 22.8717 0.8881 0.1304 17.3730 0.8086 0.2410 20.7641 0.8818 0.1380 23.3206 0.9077 0.1023 1.45
[…]
IntrinsicAnything [5] 20.7927 0.8721 0.1582 - - - - - - - - - 6.93
StableMaterial [23] 23.8992 0.9050 0.1032 18.7430 0.8321 0.2207 21.4639 0.9029 0.1168 22.8837 0.9072 0.0914 0.53
Ours w/o e2e 24.6155 0.8966 0.1102 19.8342 0.8576 0.2321 20.8168 0.8769 0.1391 24.3308 0.9125 0.0899 3.09
Ours w/o re-render 27.2149 0.9219 0.0862 22.9245 0.8474 0.1968 23.3083 0.9043 0.1122 26.7900 0.9341 0.0687 0.07
Ours 27.5034 0.9240 0.0845 23.4598 0.8598 0.1937 23.5158 0.9040 0.1114 27.1192 0.9366 0.0640 0.07
Table 1: Comparison of our method with others on the image space decomposition task. We highlight the best, second-best, and third-best results for each metric. ”Ours w/o e2e” denotes our method without the scheduler fix and perceptual loss for end-to-end training. ”Ours w/o re-render” denotes our method without the re-render loss. The first two methods are non-diffusion-based, while the middle four methods are diffusion-based.
We aim to apply SuperMat to decompose material for 3D objects with only RGB textures. Several previous works [29, 4, 45, 35] use techniques like SDS to bridge the gap between dimensions, but such methods are not applicable to our deterministic model, and suffer from slow computational speed and mode seeking issues. Therefore, we explore using backprojection to transfer 2D decomposition results from image space to UV space, enabling their application to 3D models. Notably, SuperMat is a single-view model, meaning that one inference can only decompose the visible region from a single viewpoint. To address this problem, we propose estimating materials from multiple viewpoints and integrating them. With multi-view decomposition results from SuperMat, the UV refinement network aims at inpainting missing areas, eliminating discrepancies between different views, and improving the overall quality of the generated materials.
Backprojection and UV Blending. For each viewpoint, we backproject SuperMat’s decomposition results from image space onto the model’s UV texture map, obtaining the partial albedo UV map $C_{d}^{i}$, RM UV map $C_{rm}^{i}$, and mask UV map $M^{i}$ for viewpoint $i$. We then blend the backprojection results from all viewpoints in an averaged manner. For albedo, this can be formulated as:
$\begin{split}M=\sum_{i}M^{i},\\ C_{d}=\frac{\sum_{i}C_{d}^{i}}{(M+\epsilon)}.\end{split}$ (4)
The RM follows the same computation method. The blended partial maps $C_{d}$, $C_{rm}$, and the partial mask $M$ are then input into the subsequent UV refinement one-step model.
UV Refinement One-Step Model. Due to the limited number of viewpoints, the blended partial maps often have missing areas. Additionally, slight variances may exist between SuperMat’s results from different views. To address these issues, we leverage diffusion models for UV refinement: we fine-tune Stable Diffusion into a UV refinement model, which takes a flawed partial material UV map, a mask UV map, and a position UV map as inputs, and outputs completed, variance-corrected, and refined material UV maps. Specifically, the material maps are encoded by a VAE and concatenated with downsampled mask and position maps, and then processed by UNet with an expanded 8-channel conv_in. We also modify the model’s scheduler to support single-step inference and enable end-to-end training, resulting a deterministic one-step model.
Material Decomposition Pipeline for 3D Objects. As shown in Fig. 3, with the fine-tuned UV refinement network, we propose a complete pipeline for the material decomposition of 3D objects. Given an object with only RGB texture, we first render it from six orthogonal viewpoints. The rendered results are then batch-processed by SuperMat for material decomposition. These materials, including albedo and RM, are backprojected into UV space and blended. The blended partial maps are then refined by the UV refinement one-step model and the final material UV maps are generated. This process requires only three inference steps—one for SuperMat, one for refining albedo, and one for refining RM—allowing our pipeline to complete the material decomposition of a 3D object in just 3 seconds.
4 Experiments
Figure 4: Comparison of our method with others on albedo estimation results.
Figure 5: Comparison of our method with others on metallic and roughness estimation results. We combined both materials into a single image, with the left half representing metallic and the right half showing roughness. It can be observed that these two materials exhibit more variation and are more challenging to estimate compared to albedo.
Figure 6: Results of UV refinement one-step model.
Figure 7: Qualitative ablation results for SuperMat. The first row shows albedo, while the second row shows metallic and roughness.
Figure 8: Results of decomposition for 3D objects. We show the comparison between the original and decomposed models under the same environment map, along with the relighting outcomes of new models.
Figure 9: Qualitative ablation results on UV refinement network.
Quantitative Comparison of Image Space Decomposition. We randomly select 64 unseen items each from the Objaverse [8] and BlenderVault [23] datasets. Our method is tested against 6 baseline approaches on these 128 objects. For each object, the model processes 18 inputs corresponding to 3 lighting conditions and 6 viewpoints, predicting 3 (or 1) types of materials and generating 54 (or 18) texture maps in total. All outputs are used to compute metrics, with the final averaged results shown in Tab.1. We use PSNR, SSIM, and LPIPS metrics to evaluate albedo, metallic, and roughness. Additionally, for methods that provide all material types, we compute metrics on the rendering under relighting conditions using the ground truth normals and positions and the predicted values. The results demonstrate that our approach significantly outperforms the state-of-the-art in image space decomposition.
Beyond image quality, we compare inference speeds by calculating the total time required for each method to process the test set, normalized by the dataset size and number of predicted material types. All tests are conducted with a batch size of 1. For consistency in speed comparison, StableMaterial and IID, which typically average over 10 predictions per input, are modified to perform a single prediction. The “Time(s)” column in Tab.1 reflects the time required by each method to generate a single material prediction on an NVIDIA A100 GPU with a batch size of 1. Traditional, non-diffusion-based methods like Derender3D and IIR generally show faster speeds. While StableMaterial attempts to predict multiple materials simultaneously by adding extra channels to conv_in and conv_out, its lack of structural design adjustments limits prediction quality and requires multi-step denoising, resulting in slower performance than traditional methods. In contrast, our approach achieves high-quality material decomposition with speeds comparable to traditional methods.
Qualitative Comparison of Image Space Decomposition. Fig.4 compares albedo prediction results across various cases, including artist-designed objects and real-world scans, showcasing the performance of our method against other approaches. SuperMat demonstrates superior ability in removing highlights and shadows, accurately capturing the intrinsic color, even on challenging metallic surfaces. In Fig.5, we present metallic and roughness predictions from our method alongside other baselines. SuperMat outperforms the others with well-defined boundaries between metallic and non-metallic regions and achieves consistency within material types while preserving high-frequency details. Moreover, our results closely align with physical realism, addressing a significant challenge that previous methods struggled to overcome. Results with a wider variety of inputs will be presented in the supplementary materials.
Results of UV Refinement and Material Estimation for 3D Objects. Fig.6 illustrates the effectiveness of the UV refinement one-step model in inpainting missing areas and harmonizing viewpoint variances in SuperMat’s outputs. Fig.8 displays the results from our material estimation pipeline for 3D models. The first column shows the rendering of the input models with only RGB texture, where all surfaces exhibit uniform reflectance behavior. After applying our method, the decomposed materials enable physically-based rendering with surface-specific reflectance properties, as demonstrated in the second and third columns under different lighting conditions.
Ablation Studies. In Tab.1, we present ablation studies on two core components of SuperMat: end-to-end (e2e) training and re-render loss. E2E with perceptual loss significantly enhances model performance and efficiency. In fact, within the e2e framework, we experiment with three types of loss functions, with perceptual loss achieving the best results. This suggests that direct supervision of the fully denoised predicted material maps, rather than latent noise, plays a critical role in enhancing the model’s capability for this specific task. Detailed findings are available in the supplementary materials. The addition of re-render loss, facilitated by the e2e framework, further improves prediction quality, particularly under relighting conditions. Qualitative results of these ablations are shown in Fig. 7. The ‘w/o e2e’ setting, where the model is trained without perceptual and re-render losses, limits its capacity to remove lighting effects effectively. In contrast, training ‘w/o re-render’ loss leads to minor inaccuracies in fine material details. Note that in the ’w/o e2e’ setting, there are slight differences in the implementation of the UNet. Detailed explanations will be provided in the supplementary materials.
For the decomposition pipeline for 3D objects, we conduct additional ablations on the UV refinement network. Due to limitations from only six orthogonal views, achieving complete texture coverage, accurate merging of overlapping regions, and consistent material transitions are challenging. As shown in Fig.9, omitting the UV refinement leads to incomplete textures, visible color-separation lines, and unnatural transitions between materials, underscoring the importance of this component.
5 Conclusions
In this paper, we present SuperMat, an innovative image space material decomposition one-step model fine-tuned from Stable Diffusion with task-specific losses in an end-to-end framework. SuperMat enables simultaneous multi-material prediction and supports single-step inference, achieving state-of-the-art quality and outperforming previous diffusion-based methods in efficiency. Through the proposed UV refinement network, we extend SuperMat to decomposition tasks for 3D objects, where it demonstrates exceptional performance and applicability.
Supplementary Material
6 Additional Details
Implementation Details. We report some implementation details as follows: 1) In the UV refinement one-step model, the input channels are expanded to 8, with the weights of the additional 4 channels initialized to 0. 2) In the ablation experiment, the structure of SuperMat under the “w/o e2e” setting is slightly different. Without single-step inference, we can only train the diffusion model in a denoising task manner, meaning that during the single-step denoising process, the latents that the UNet receives are not encoded from a clean rendered image, but rather a noisy albedo and noisy RM. Therefore, on top of SuperMat, we additionally replicate the conv_in and the first DownBlock as independent parts of structural expert branches to map the inputs from two different domains to similar distributions. These are then fused in an averaged manner before being fed into the shared modules. 3) In the re-render loss implementation, each time we perform relighting, we randomly select a lighting condition from a set of 716 environment maps, covering nearly all possible lighting scenarios.
Training Details. Both models are fine-tuned from Stable Diffusion 2.1. We train SuperMat using the AdamW optimizer with a learning rate of $51e-$ 5 on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 16, for a total of 115 epochs. The UV refinement one-step model is trained with the AdamW optimizer at a learning rate of $52e-$ 5, also on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 16, for 40 epochs. The images are resized to resolutions of $512\times 512$ and $1024\times 1024$, used for SuperMat and the UV refinement one-step model, respectively.
Training Datasets. In this paper, we collect 3D objects from Objaverse [8]. For each object, we render material maps, normal maps, and position maps from six orthogonal viewpoints and generate PBR results under three different environment lighting conditions from the same viewpoints. We then filter high-quality data using methods such as DINO score considerations [28], resulting in a dataset of 100k objects for the image space decomposition task. Following the approach outlined in Sec. 3.3, we further refine this data for the UV refinement network by selecting cases where the missing area in blended partial UV maps is under $50\%$. This process results in a final dataset of 40k objects for training the refinement one-step model.
[…]
7 Additional Visualizations
In Figures 10111213, we present additional results of SuperMat on the Objaverse and BlenderVault test datasets, covering both artist-designed and real-world scanned objects. Fig. 14 highlights more decomposition results of real-world objects extracted from captured images. Finally, Fig. 15 illustrates more decomposition results on 3D objects.
Figure 10: Additional albedo comparison from the Objaverse test dataset.
Figure 11: Additional albedo comparison from the BlenderVault test dataset.
Figure 12: Additional metallic and roughness comparison from the Objaverse test dataset. The metallic maps are shown on the left side (M), while the roughness maps are shown on the right side (R).
Figure 13: Additional metallic and roughness comparison from the BlenderVault test dataset. The metallic maps are shown on the left side (M), while the roughness maps are shown on the right side (R).
Figure 14: Additional results on real-world objects. The objects are extracted from captured images by SAM [19].
Albedo Metallic Roughness Relighting
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
L1 26.3745 0.9159 0.0987 21.8087 0.8733 0.2113 22.4262 0.9013 0.1308 25.8089 0.9300 0.0765
SSI 23.8053 0.9073 0.1067 21.6018 0.8378 0.2097 22.8090 0.9063 0.1261 23.8945 0.9242 0.0813
Perceptual 27.7214 0.9252 0.0771 23.1332 0.8799 0.1584 24.7161 0.9154 0.0926 28.0626 0.9424 0.0548
Table 2: Quantitative comparison across different loss functions. We highlight the best results for each metric.
Figure 15: Additional results of decomposition for 3D objects.
8 Additional Experiments
We show the detailed quantitative comparison in Tables 345 and  6. Additionally, without re-render loss, we experiment with three types of loss functions in the end-to-end framework: L1 loss, shift and scale invariant (SSI) loss [31], and perceptual loss. The performance of models trained with these loss designs is evaluated on the same test dataset, and the results are shown in Tab. 2. Among these losses, the perceptual loss, which is ultimately adopted by SuperMat, demonstrates the best performance.
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
Derender3D 0.0094 22.0689 0.8529 0.1754 0.0104 22.4523 0.8919 0.1489
IIR 0.0130 21.3350 0.8705 0.1587 0.0202 20.5980 0.8654 0.1672
IID 0.0081 23.3557 0.8916 0.1212 0.0122 22.3877 0.8845 0.1395
[…]
IntrinsicAnything 0.0121 21.3502 0.8743 0.1481 0.0219 20.2352 0.8700 0.1682
StableMaterial 0.0101 23.8964 0.9038 0.0985 0.0143 23.9020 0.9061 0.1079
Ours w/o e2e 0.0058 24.9261 0.8997 0.0986 0.0102 24.3049 0.8935 0.1219
Ours w/o re-render 0.0023 27.7214 0.9252 0.0771 0.0043 26.7084 0.9185 0.0953
Ours 0.0020 28.1156 0.9284 0.0753 0.0044 26.8911 0.9195 0.0938
Table 3: Quantitative comparison on albedo. We highlight the best, second-best, and third-best results for each metric.
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0471 16.1588 0.8401 0.1997 0.0515 17.6046 0.7835 0.2732
IID 0.0343 17.3157 0.8430 0.1954 0.0431 17.4303 0.7742 0.2866
[…]
StableMaterial 0.0450 16.6922 0.8487 0.2034 0.0515 20.7938 0.8156 0.2379
Ours w/o e2e 0.0353 18.7623 0.8696 0.1879 0.0470 20.9061 0.8455 0.2764
Ours w/o re-render 0.0134 23.1332 0.8799 0.1584 0.0330 22.7159 0.8150 0.2352
Ours 0.0125 23.5208 0.8845 0.1573 0.0291 23.3987 0.8351 0.2302
Table 4: Quantitative comparison on metallic. We highlight the best, second-best, and third-best results for each metric.
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0195 20.7515 0.8877 0.1367 0.0334 19.1632 0.8621 0.1790
IID 0.0176 21.5627 0.8917 0.1174 0.0295 19.9655 0.8718 0.1586
[…]
StableMaterial 0.0133 21.9898 0.9096 0.1024 0.0251 20.9379 0.8962 0.1312
Ours w/o e2e 0.0161 21.1543 0.8858 0.1148 0.0278 20.4794 0.8679 0.1633
Ours w/o re-render 0.0070 24.7161 0.9154 0.0926 0.0221 21.9006 0.8933 0.1317
Ours 0.0070 24.7642 0.9144 0.0915 0.0214 22.2674 0.8935 0.1313
Table 5: Quantitative comparison on roughness. We highlight the best, second-best, and third-best results for each metric.
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0124 22.5646 0.9021 0.1116 0.0296 20.4861 0.8782 0.1441
IID 0.0085 24.3354 0.9144 0.0903 0.0160 22.3058 0.9009 0.1142
[…]
StableMaterial 0.0115 23.5445 0.9121 0.0814 0.0203 22.2230 0.9022 0.1013
Ours w/o e2e 0.0072 25.5109 0.9219 0.0739 0.0171 23.1506 0.9031 0.1059
Ours w/o re-render 0.0041 28.0626 0.9424 0.0548 0.0123 25.5175 0.9259 0.0826
Ours 0.0035 28.4900 0.9465 0.0510 0.0106 25.7484 0.9266 0.0770
Table 6: Quantitative comparison on relighting. We highlight the best, second-best, and third-best results for each metric.